OpenStudy (anonymous):

Colleen's car is depreciating at a rate of 9% per year. she paid $24,500 for it in 2010 a) write a function to model the value of the car in t years

OpenStudy (johnweldon1993):

so we have V = Vi (1 - r)^t where v is value Vi is initial value r is rate and t is time in years V = 24500(1-.09)^t V = 24500(.91)^t

OpenStudy (anonymous):

Ohh so its like when its increasing, but a 1- instead of a 1+?

OpenStudy (johnweldon1993):

correct....since this would be exponential decay...it is 1 - the rate....instead of exponential growth being 1 + rate

OpenStudy (anonymous):

Thanks! Do you mind helping me in another question; ill post it

OpenStudy (johnweldon1993):

yeah sure...if i can that is